OpenStudy (anonymous):

Solve this Proportion: 3/2y = 16/y+2 Solve this Proportion: 3/2y = 16/y+2 @Mathematics

OpenStudy (anonymous):

Multiply through by 2y

OpenStudy (anonymous):

3y+6 = 32y then 29y =6 - > y =6/29

OpenStudy (anonymous):

Were did you get 6?

OpenStudy (anonymous):

3*(y+2) = 16*2y

OpenStudy (anonymous):

Ok thank you! :D